OpenStudy (anonymous):

Related Rates! A ladder 16 feet long is learning against the wall of a building. The base of the ladder is moving away from the wall at a rate of 1 foot per second. Find the rate at which the angle theta between the ladder and the wall of the building is changing when the base of the ladder is 8 feet from the wall.

OpenStudy (anonymous):

I've already drawn a picture and come up with the relationship: sin(theta) = x/16. Is that correct?

OpenStudy (anonymous):

you re right

razor99 (razor99):

what is that @ carabell

OpenStudy (anonymous):

Then here's my issue: If you try to differentiate that--you know, to get the rate of change--you don't end up using the x. sin(theta) = x/16 --> cos(theta) = 1/16 dx/dt. dx/dt is just 1. And the x=8 doesn't come into play. Am I making sense? I'm just concerned.

OpenStudy (anonymous):

Oh, oops. There's probably a dtheta/dt in there as well...

OpenStudy (anonymous):

What do you get if you differentiate in respect to time? Or do you have to do that at all?

OpenStudy (anonymous):

rate = d(theta)/dt = I/(16 x Cos theta)

OpenStudy (anonymous):

Is is possible for you explain how you got that?

OpenStudy (anonymous):

cos(theta) d(theta)/d (time) = 1/16 dx/dt so from question, we know that dx/dt = 1 foot per sec hence cos(theta) d(theta)/d (time) = 1/16 (1) d(theta)/dt = I/(16 x Cos theta)

OpenStudy (anonymous):

Oh! And that x is multiplication?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What I don't understand is why do they, in the question, specify x=8 if you're not going to use it?

OpenStudy (anonymous):

may be to confuse..... you just need to get the idea....

OpenStudy (anonymous):

Okay. Thank you for your help.

OpenStudy (anonymous):

if you ll use d(8+x)/dt , still then you ll get dx/dt because 8 is constant

OpenStudy (anonymous):

pleasure.....